Pure acid is to be added to a 10% acid solution to obtain 90l of 56% solution. What amounts should be used?

Answer by stanbon(75887)

Pure acid is to be added to a 10% acid solution to obtain 90l of 56% solution. What amounts should be used?
-------
Equation:
acid + acid = acid
x + 0.10(90L-x) = 0.56*90
------
100x + 10*90L - 10x = 56*90
----
90x = 46*90
x = 46 L (amt. of pure acid to add)
==================
Cheers,
Stan H.
===================